OpenStudy (anonymous):

convert to radians 25^0 medall!!!!!!!!!!

OpenStudy (anonymous):

@Hero

Nnesha (nnesha):

is it 25 to the 0 power ? \[25^0\] like this ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

ohh it's a degree not 0 :3

OpenStudy (anonymous):

ow lol sorry

Nnesha (nnesha):

25 degree not 0 anything to the 0 power is just equal to 1 (Nnesha)^0 = 1 ( x)^0 = 1 45^0 = 1

OpenStudy (anonymous):

so its 1 right:)

Nnesha (nnesha):

nope iit's not 25^0 it's 25 degree

Nnesha (nnesha):

 like this

OpenStudy (anonymous):

but it says to convert to radians 25 degrees

Nnesha (nnesha):

formula to convert degree to radians is \[\huge\rm d \times \frac{ \pi }{ 180}\] d=degree

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

im lost sorry

OpenStudy (anonymous):

so the answer is 25 degrees?

OpenStudy (anonymous):

or 5pi/36?

Nnesha (nnesha):

nope Questions is convert 25 degree to radians i asked u is it 0 and ur answer was yes but no it's not 0 it's 25 degree now use that formula to convert it to radians

Nnesha (nnesha):

yes that's right

OpenStudy (anonymous):

thanks and that the answer right:)

Nnesha (nnesha):

yes 5pi/36 is right :-)

OpenStudy (anonymous):

thanks so much!!!!

Nnesha (nnesha):

np